# Universal Biological Sequence Reranking for Improved De Novo Peptide Sequencing

### v1

@@ 1 Introduction
Over the past two decades, de novo sequencing has made substantial progress, evolving from graph-theoretic and dynamic programming-based methods to more sophisticated approaches driven by deep learning (Ma et al., 2003; Liu et al., 2022; LeCun et al., 2015; Zhang et al., 2025a; Gao et al., 2023). DeepNovo (Tran et al., 2017) was the first to apply deep learning to de novo sequencing, which inspired a series of subsequent models(Zhou et al., 2017; Karunratanakul et al., 2019; Yang et al., 2019; Liu et al., 2023). More recently, Transformer architectures have been introduced to model de novo sequencing as a machine translation task (Yilmaz et al., 2022; Zhang et al., 2024b; Mao et al., 2023; Eloff et al., 2023a; Yang et al., 2024b; Xia et al., 2024a). Building upon this foundation, ContraNovo (Jin et al., 2024) further advanced the field by incorporating multimodal alignment strategies, achieving state-of-the-art performance.
@@ 2.1 De Novo Peptide Sequencing
De novo sequencing algorithms have witnessed the adaption from early dynamic programming with rule-based scoring functions (Ma et al., 2003; Chi et al., 2010; Ma, 2015) to deep-learning based end-to-end regression, mainly based on transformer architecture (Yilmaz et al., 2022; Mao et al., 2023; Eloff et al., 2023a; Zhang et al., 2024a; Yang et al., 2024b; Jin et al., 2024; Zhou et al., 2024; Eloff et al., 2023b; Xia et al., 2024a; Zhang et al., 2025c; Yang et al., 2024a; Xia et al., 2024b).
@@ 2.2 Candidate Reranking
In the reranking task, methods are typically categorized into three types: point-wise, pair-wise, and list-wise (Zhuang et al., 2023). The point-wise method independently evaluates the relevance of a single query-candidate pair (Nogueira et al., 2019, 2020). The pair-wise method assesses the relative relevance between two candidate pairs for a given query (Burges et al., 2005; Burges, 2010; Ouyang et al., 2022; Yin et al., 2023; Jiang et al., 2023; Wei et al., 2025). The list-wise method considers the relevance of all candidate pairs for each query collectively, utilizing all candidate features, which enhances performance potential (Han et al., 2020; Gao et al., 2021; Ren et al., 2021; Cao et al., 2007; Wei et al., 2023; Xia et al., 2008; Zhang et al., 2025b).
@@ 3.2 Spectrum and Peptide Embedding
Peptide Candidate Embedding De novo sequencing is a mass-centric task. Therefore, the prefix and suffix masses of a residual are also embedded in addition to a learnable amino acid embedding, following ContraNovo. Given the model dimension $d$, the dimensions of the learnable embedding, prefix, and suffix, denoted as $d_{\textnormal{res}},d_{\textnormal{prefix}},d_{\textnormal{suffix}}$, are set to $\frac{d}{2},\frac{d}{4}$ and $\frac{d}{4}$ respectively. Moreover, the precursor of spectrum are embedded into $d_{\textnormal{prec}}$ ($\frac{d}{2}$) dimensions as the sum of $d_{\textnormal{prec}}$-dimensional precursor mass vector and precursor charge vector (Jin et al., 2024). All masses $m$ are embedded using fixed sinusoidal positional embedding:
@@ 4.1 Experiment Setup
Datasets. Following the precedent set by recent studies (Yilmaz et al., 2023; Zhang et al., 2024c), we employ three public peptide-spectrum match (PSMs) datasets: MassIVE-KB (Wang et al., 2018) for training, and 9-species-V1 (Tran et al., 2017) and 9-species-V2 (Yilmaz et al., 2023) for evaluation, enabling comparisons with state-of-the-art de novo peptide sequencing methods. Detailed dataset information is provided in the Appendix A.

### v2

@@ 1 Introduction
Over the past two decades, de novo sequencing has made substantial progress, evolving from graph-theoretic and dynamic programming-based methods to more sophisticated approaches driven by deep learning (Ma et al., 2003; LeCun et al., 2015; Gao et al., 2023). DeepNovo (Tran et al., 2017) was the first to apply deep learning to de novo sequencing, which inspired a series of subsequent models(Zhou et al., 2017; Karunratanakul et al., 2019; Yang et al., 2019; Liu et al., 2023). More recently, Transformer architectures have been introduced to model de novo sequencing as a machine translation task (Yilmaz et al., 2022; Mao et al., 2023; Eloff et al., 2023a; Yang et al., 2024b; Xia et al., 2024a). Building upon this foundation, ContraNovo (Jin et al., 2024) further advanced the field by incorporating multimodal alignment strategies, achieving state-of-the-art performance.
@@ 2.1 De Novo Peptide Sequencing
De novo sequencing algorithms have witnessed the adaption from early dynamic programming with rule-based scoring functions (Ma et al., 2003; Chi et al., 2010; Ma, 2015) to deep-learning based end-to-end regression, mainly based on transformer architecture (Yilmaz et al., 2022; Mao et al., 2023; Eloff et al., 2023a; Yang et al., 2024b; Jin et al., 2024; Zhou et al., 2024; Eloff et al., 2023b; Xia et al., 2024a; Zhang et al., 2025; Yang et al., 2024a; Xia et al., 2024b).
@@ 2.2 Candidate Reranking
In the reranking task, methods are typically categorized into three types: point-wise, pair-wise, and list-wise (Zhuang et al., 2023). The point-wise method independently evaluates the relevance of a single query-candidate pair (Nogueira et al., 2019, 2020). The pair-wise method assesses the relative relevance between two candidate pairs for a given query (Burges et al., 2005; Burges, 2010; Ouyang et al., 2022; Jiang et al., 2023). The list-wise method considers the relevance of all candidate pairs for each query collectively, utilizing all candidate features, which enhances performance potential (Han et al., 2020; Gao et al., 2021; Ren et al., 2021; Cao et al., 2007; Xia et al., 2008).
@@ 3.2 Spectrum and Peptide Embedding
Peptide Candidate Embedding De novo sequencing is a mass-centric task. Therefore, the prefix and suffix masses of a residual are also embedded in addition to a learnable amino acid embedding, following ContraNovo. Given the model dimension $d$, the dimensions of the learnable embedding, prefix, and suffix, denoted as $d_{\textnormal{res}},d_{\textnormal{prefix}},d_{\textnormal{suffix}}$, are set to $\frac{d}{2},\frac{d}{4}$ and $\frac{d}{4}$ respectively. Moreover, the precursor of spectrum are embedded into $d_{\textnormal{prec}}$ ($\frac{d}{2}$) dimensions as the sum of $d_{\textnormal{prec}}$-dimensional precursor mass vector and precursor charge vector (Jin et al., 2023). All masses $m$ are embedded using fixed sinusoidal positional embedding:
@@ 4.1 Experiment Setup
Datasets. Following the precedent set by recent studies (Yilmaz et al., 2023; Zhang et al., 2024), we employ three public peptide-spectrum match (PSMs) datasets: MassIVE-KB (Wang et al., 2018) for training, and 9-species-V1 (Tran et al., 2017) and 9-species-V2 (Yilmaz et al., 2023) for evaluation, enabling comparisons with state-of-the-art de novo peptide sequencing methods. Detailed dataset information is provided in the Appendix A.